OpenStudy (anonymous):

how do you isolate i^2 from (64/n) - (64i^2)/n^3?

OpenStudy (jdoe0001):

maybe it's just me, but there's no "equation" per se, just an expression, thus you can't quite isolate it

OpenStudy (jdoe0001):

not to mention that \(i^2\) is a constant \(i^2 = (-1)(-1) = -1\)

OpenStudy (anonymous):

I think I should post the original problem. ^^ thank you though.

OpenStudy (anonymous):

\[\frac{ 64 }{ n }-\frac{ 64 i^2 }{ n^3 } = 0\]

OpenStudy (anonymous):

is this the equation?